I need help solving a few questions on Chapter 4.3 Solving Systems of Linear Equations by the Substitution Method. Question 1 How do I solve this problem? 2x-3y=11
-4x+3y=-19

Answer by Alan3354(69443)

How do I solve this problem?
+2x-3y=11
-4x+3y=-19
------------ Add them